# Reducing Numerical Precision Requirements in Quantum Chemistry Calculations

### v1

@@ IV.6 Mixed Precision Purification
Finally, we will revisit the result of Sec. IV.1 regarding the precision needed for multiplying matrices and for accumulation. The findings of Fig. 5 indicate that the matrices used in purification may be stored in FP32 as long as the multiplication upcasts them to FP64. This could be utilized in current libraries that implement purification [60, 61, 33, 62, 63] to reduce communication and data transfer costs, even without considering the Ozaki scheme. To validate this finding, we modify the NTPoly code [33], which implements the TRS4 method using sparse matrix algebra. During the multiplication process, we down cast all input matrix elements back and forth from FP32 to simulate the loss of precision. As a test matrix, we use the Hamiltonians coming from a BigDFT calculation of a bulk silicon supercell (3240 atoms) and an NTChem calculation of a water cluster (573 molecules).
@@
V Conclusion
In this work, we have examined the specific floating point precision requirements of a representative kernel from quantum chemistry calculations: calculation of the single particle density matrix using density matrix purification. Exploiting MPFR to emulate arbitrary mantissa sizes, we found that double precisions affords an unnecessarily high level of precision. If a precision between single and double precision is used (for example, with an effective 37 bit mantissa), a reliable result can be obtained. We further identified that single precision is sufficient if accumulation is done in a higher precision; libraries that implement density matrix purification or similar algorithms may immediately exploit this fact to reduce communication and data transfer costs. Conversely, in the case of calculations with large basis sets, the exponent range of double precision can limit the quality of the final result.
To further take advantage of this reduced precision requirement, we proposed the use of the Ozaki scheme with a smaller number of splits. We found that the reduced precision requirements of purification leads to a reduction in the number of multiplications needed for the Ozaki scheme by about a factor of two. In this work, we have only examined the reliability of this approach, and not implemented an optimized version. Nonetheless, on an RTX A6000 GPU the Ozaki-scheme has already been demonstrated to outperform standard double precision multiplication calls [45]. Furthermore, on an A100 a similar algorithm to purification was shown to require less time to solution then dense diagonalization [37]. Hence, a combination of the two may provide practical benefit in the short term (in particular, a node parallel version). More importantly, if future architectures are developed with an even higher ratio of low precision multiplication to FP64, our work shows that a performance improvement could be realized without sacrificing precision.
The methodology developed here may be straightforwardly applied to a number of other matrix multiplication based algorithms in quantum chemistry (particularly many-body methods like MP2 and Coupled Cluster). Software emulation of low precision results can substantially increase the time to solution, however many practical problems remain in reach, particularly with the development of appropriate libraries for multi-precision linear algebra [64]. Another approach may be to use tools like veritracer to instrument electronic structure codes and measure floating point errors [65, 66].

### v2

@@ 4.6 Mixed Precision Purification
Finally, we will revisit the result of Sec. 4.1 regarding the precision needed for multiplying matrices and for accumulation. The findings of Fig. 5 indicate that the matrices used in purification may be stored in FP32 as long as the multiplication upcasts them to FP64. This could be utilized in current libraries that implement purification 75, 76, 39, 77, 78 to reduce communication and data transfer costs, even without considering the Ozaki scheme. To validate this finding, we modify the NTPoly code 39, which implements the TRS4 method using sparse matrix algebra. During the multiplication process, we down cast all input matrix elements back and forth from FP32 to simulate the loss of precision. As a test matrix, we use the Hamiltonians coming from a BigDFT calculation of a bulk silicon supercell (3240 atoms) and an NTChem calculation of a water cluster (573 molecules).
@@
5 Conclusion
In this work, we have examined the specific floating point precision requirements of a representative kernel from quantum chemistry calculations: calculation of the single particle density matrix using density matrix purification. Exploiting MPFR to emulate arbitrary mantissa sizes, we found that double precision affords an unnecessarily high level of precision. If a precision between single and double precision is used (for example, with an effective 37 bit mantissa), a reliable result can be obtained. We further identified that single precision is sufficient for many practical problems if accumulation is done in a higher precision; libraries that implement density matrix purification or similar algorithms may immediately exploit this fact to reduce communication and data transfer costs.
To further take advantage of this reduced precision requirement, we proposed the use of the Ozaki scheme with a smaller number of splits. We found that the reduced precision requirements of purification leads to a reduction in the number of multiplications needed for the Ozaki scheme by about a factor of two. In this work, we have only examined the reliability of this approach, and not implemented an optimized version. Nonetheless, on an RTX A6000 GPU the Ozaki scheme has already been demonstrated to outperform standard double precision multiplication calls 57. Furthermore, on an A100 a similar algorithm to purification was shown to require less time to solution then dense diagonalization 43. Hence, a combination of the two may provide practical benefit in the short term (in particular, a node parallel version). More importantly, if future architectures are developed with an even higher ratio of low precision multiplication to FP64, our work shows that a performance improvement could be realized without sacrificing precision.
The methodology developed here may be straightforwardly applied to a number of other matrix multiplication based algorithms in quantum chemistry — particularly many-body methods like MP2 and Coupled Cluster. For materials science applications based on planewaves, finite elements, etc., the Ozaki scheme may be used to accelerate steps like orbital orthogonalization or subspace diagonalization. Software emulation of low precision results can substantially increase the time to solution, however many practical problems remain in reach, particularly with the development of appropriate libraries for multi-precision linear algebra 79. Another approach may be to introduce artificial numerical noise into codes 80 or use tools like veritracer to automatically instrument electronic structure codes and measure floating point errors 81, 82.